# Details Enhancement in Unsigned Distance Field Learning for High-fidelity 3D Surface Reconstruction

### v1

@@ 4.3 Ablation studies
We assess the impact of using an unconditioned SIREN network on the performance of our method by comparing it to versions of the SIREN network that utilize absolute value and softplus function in the output layer, respectively. As shown in Table 3, the SIREN network with an absolute value output tends to learn a “fake” UDF, which behaves similarly to an SDF in modeling watertight models. On the other hand, due to the vanishing gradient effect of softplus, the reconstructed mesh using the SIREN network with a softplus output is typically over-smoothed.
Normal alignment
To evaluate the contribution of normal alignment to reconstruct geometric details in our model, we performance comparisons between versions of our method with and without the normal alignment loss. The results highlighting the improvements are documented in Table 3.
Table 3: Ablation studies on the model “Stonewall” in the Stanford 3D Scene Dataset.
CD-mean CD-median $F1^{0.01}$ $F1^{0.005}$
SIREN+Abs. 5.27 3.07 87.65 77.87
SIREN+softplus 2.91 2.71 99.79 91.08
w/o weighted Eikonal 2.87 2.54 99.26 90.98
w/o normal alignment 3.20 2.97 99.36 88.68
DEUDF 2.69 2.51 99.95 93.52
Weighted Eikonal
We explore the effects of different configurations of the Eikonal loss. Specifically, we compare our method using a standard Eikonal loss that is applied uniformly to all sample points with our adaptively weighted Eikonal loss. We observe that the standard Eikonal loss results in learned UDFs with lower accuracy near the zero level sets, leading to numerous small holes in the extracted meshes. In contrast, our adaptively weighted Eikonal loss more effectively address the issues of vanishing gradient, and stabilizes the learning process, thereby yielding meshes with higher quality. Due to the uniform Eikonal constraint obstructing the UDF learning process at the zero level set, slightly higher distance values are observed around the target surface. To accommodate these inaccuracies, we applied a higher threshold of 0.01, instead of the default value of 0.0025 used in DCUDF, for extracting the surface.
@@
Figure 4: Visual results from the ablation studies: (a) Applying the absolute value to the output of the SIREN network. (b) Applying the softplus function to the output of the SIREN network. (c) Using uniform Eikonal constraints. (d) Removing normal alignment. (f) Applying all components.
5 Conclusions
This paper presents an improved UDF learning method for high fidelity 3D surface reconstruction. The method integrates novel UDF representation without absolute or softplus, normal alignment, adaptively weighted Eikonal constraint and SIREN network to learn more accurate UDFs. Our DEUDF can learn not only geometry details but also geometry boundaries, thereby maintaining better topology. Extensive experiments illustrate that our method produces low Chamfer distances and better topology outperforming state-of-the-art methods.
Since our method needs normals as guidance, the input point cloud should be dense. Moreover, to learn model details also needs dense input. In the future, we will investigate to learn uneven density point cloud, where the points are sparse in smooth regions and dense in detailed regions. Another limitation is highly noisy point clouds, which are challenge for detailed model learning. We will explore how to learn noise inputs in the future.
Appendix A Appendix
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(a) CAP-UDF
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(b) DUDF
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@ Appendix A Appendix
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
We present additional comparisons with CAP-UDF [3] and LevelSetUDF [4] by their original implementations as illustrated in Figure 5 and detailed in Table 4. The accuracy is lower than using DCUDF [8] for surface extraction. Our method produces higher quality results, characterized by enhanced geometric details and smoother shape boundaries. In Figure 6 and 7, we show more results for models with details and models with complex topology and inner structures. The surfaces are all extracted by DCUDF except DUDF [17] by the original method. Our method outperforms DUDF, CAP-UDF and LevelSetUDF in terms of accuracy and topology.
Table 4: Quantitative comparisons with the original results of CAP-UDF and LevelSetUDF. Chamfer distances are measured in the unit of ×$\times 10^{-3}$. We randomly sampled 100K points from each ground truth mesh and then computed both CDs and F-scores as reference performance metrics. Therefore, the closer the actual performance metrics are to these references, the higher the quality of the results produced by the method.
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 7: More visual comparisons with DUDF, CAPUDF and LevelSetUDF on car models from the ShapeNet-Cars dataset showcasing better reconstruction for complex inner structures.

### v2

@@ Introduction
Table 2: Quantitative results on Stanford 3D Scene, Stanford 3D Scans (watertight) and ShapeNet-Cars. Chamfer distances are measured in the unit of ×$\times 10^{-3}$. For CAP-UDF and LevelSetUDF, their results often include small isolated components. To compute the Chamfer distance, we cleaned their meshes by removing components that were distant from the ground truth point cloud after mesh extraction using DCUDF.
Surface extraction
After obtaining the UDFs, we proceed to extract the surface from the learned UDFs. Due to the relaxation of the non-negative constraints, the target geometry does not align precisely with the zero level set. Instead, we identify the target surface as the local minimal distance surface (which can be either positive or negative) near the zero values.
One possible method for extracting the target geometry from UDFs involves explicitly using the UDF gradient, such as MeshUDF (Guillard, Stella, and Fua 2022) and GeoUDF (Ren et al. 2023), both of which are variants of the standard Marching Cubes algorithm. On each cube edge, if the gradient directions of the UDF at the two endpoints are opposite and their UDF values are below a specified threshold, a zero crossing is marked on that edge. However, this approach is not suitable for our purpose because our learned UDFs do not ensure the positivity of the distance value, and the distance values on the surface may exceed the specified UDF threshold. If this occurs, the iso-surfacing method would exclude these cubes, leading to the creation of extracted surfaces with undesired holes.
To tackle this challenge, we adopt the optimization-based method DCUDF (Hou et al. 2023), which initiates by extracting a double cover using the Marching Cubes algorithm at a small positive iso-value on the UDF. Subsequently, it shrinks the double cover to the local minimal distance surface. This method does not require the UDF to be strictly positive nor does it depend on a threshold to select candidate cubes. As a result, it effectively identifies the local minimal distance surface, yielding a high-quality triangle mesh that accurately represents the target surface.
Experiments
Refer to captionRefer to captionRefer to captionRefer to caption
(a) DUDF
Refer to captionRefer to captionRefer to captionRefer to caption
(b) LevelSetUDF
Refer to captionRefer to captionRefer to captionRefer to caption
(c) NSH
Refer to captionRefer to captionRefer to captionRefer to caption
(d) Ours
Refer to captionRefer to captionRefer to captionRefer to caption
(e) GT
Figure 2: Visual comparison with two recent UDF learning approaches, DUDF (Fainstein, Siless, and Iarussi 2024) and LevelSetUDF (Zhou et al. 2023), and one recent SDF learning method, NSH (Wang et al. 2023b), on surfaces with fine geometric details. Our method yields visually pleasing results, reconstructing significantly more details than the other methods.
Datasets
We evaluate our method using three datasets: ShapeNet-Cars (Chang et al. 2015) with 108 models [^2] We select all models whose names start with “1”., the Stanford 3D Scene Dataset (Zhou and Koltun 2013) with 5 models and the Stanford 3D Scan Repository  [^3] https://graphics.stanford.edu/data/3Dscanrep/ with 8 models. For each shape, we randomly sample 300K points as input. After learning the UDFs, we employ DCUDF with a resolution of $512^{3}$ to extract the target surface. To evaluate the accuracy, We use Chamfer distance (CD) and F-score as quantitative measures. For F-score, we set the thresholds to $0.5\%$ and $0.25\%$. Following previous methods (Zhou et al. 2022, 2023), we randomly sample 100K points from both the reconstructed surfaces and the ground truth meshes for computing CD and F-score. All points and meshes are normalized to bounding boxes whose longest edges are 2. We test on an NVIDIA Tesla V100 GPU with 32GB memory (about 5GB used for a UDF learning). It takes about 30 minutes to learn a UDF.
Results & comparisons
We compare our method with three state-of-the-art UDF learning methods: LevelSetUDF (Zhou et al. 2023), CAP-UDF (Zhou et al. 2022) and DUDF (Fainstein, Siless, and Iarussi 2024). Since we adopt DCUDF (Hou et al. 2023) for surface extraction, we also test DCUDF for the three baselines to ensure fairness of comparisons. For CAP-UDF and LevelSetUDF, we observe that DCUDF could produce better results than their original implementations in terms of Chamfer distances and visual effects. But for DUDF, the results of DCUDF are not as good as the original results. Therefore, to report the best results of the baseline methods, we choose to adopt DCUDF for extracting the zero level set from the UDF outputs from both CAP-UDF and LevelSetUDF. While DUDF uses their original results for comparisons. For the ShapeNet-Cars and Stanford 3D Scene datasets, the outputs of DCUDF remain as double-layered meshes, bypassing the min-cut based double-layer segmentation post-processing. In addition, we assess our approach against IDF (Wang, Rahmann, and Sorkine-Hornung 2022) and NSH (Wang et al. 2023b), two state-of-the-art SDF learning methods, on the watertight surfaces with fine geometric details from the Stanford 3D Scan Repository. The results are illustrated in Table 2 and Figure 2.
3D objects with fine geometric details
3D objects with complex inner structures
We further explore our method for representing 3D objects and scenes with complex inner structures. We evaluate our method on Stanford 3D Scene and ShapeNet-Cars datasets. As shown in Table 2 and Figure 4, our method is more stable on complex structures and performs optimally for keeping open boundaries.
Gaussian Noise-N(0,0.0025) Gaussian Noise-N(0,0.005)
Chamfer-L1 ()($\downarrow$)( ↓ ) F-score ()($\uparrow$)( ↑ ) Chamfer-L1 ()($\downarrow$)( ↓ ) F-score ()($\uparrow$)( ↑ )
Method Mean Median $F1^{0.005}$ $F1^{0.0025}$ Mean Median $F1^{0.005}$ $F1^{0.0025}$
CAP-UDF 5.76 5.49 92.68 43.39 6.26 5.99 87.82 37.54
LeverSetUDF 5.63 5.37 93.39 45.60 6.42 6.18 86.45 36.05
DUDF 4.27 3.83 97.91 70.57 4.61 4.11 96.25 65.04
Ours 4.07 3.79 98.74 71.89 5.64 5.35 92.74 46.18
Table 3: Quantitative results for reconstruction on noisy point clouds from Stanford 3D Scenes and 3D Scans. We add Gaussian noise $N(0,0.0025)$ and $N(0,0.005)$ to the input point clouds and use the SIREN with a frequency of 30 to resist noises.
Noisy point clouds
As illustrated in Table 3, we add $N(0,0.0025)$ and $N(0,0.005)$ Gaussian noise to the normalized point clouds. To improve the noise resistance, we reduce the SIREN frequency to 30. Our method is best at resisting small noises. DUDF is the best for large noise because it always oversmoothes the surface. However, for the same reason, DUDF is the worst at reconstructing details. Apart from DUDF, our DEUDF is better than others for large noisy data. Meanwhile, DEUDF is the best at reconstructing details.
Deviation from zero
It is important to note that the distances of the learned UDF of a surface are unlikely to be exactly zero due to inherent learning errors. Interestingly, even though we relax the non-negativity constraint in our DEUDF, the deviation from zero in our results is actually smaller than that of UDFs learned by other methods enforcing strict non-negativity. We evaluate the average absolute distance values of the reconstructed mesh points in the learned UDF to evaluate the deviation. We test on the Stanford 3D Scene and 3D Scan Repository datasets. The average deviations are as follows: DUDF: $2.60\times 10^{-3}$, CAP-UDF: $2.63\times 10^{-4}$, LevelSetUDF: $1.39\times 10^{-4}$, and our DEUDF: $1.34\times 10^{-4}$.
Ablation studies
We conduct ablation studies to demonstrate the effectiveness of each component within our method.
@@
We assess the impact of using an unconditioned SIREN network on the performance of our method by comparing it to other versions of the SIREN network that utilize absolute value and softplus function in the output layer, respectively. As shown in Figure 3 and Table 4, the absolute output layer results in large fitting error. On the other hand, due to the vanishing gradient effect of softplus, the reconstructed mesh using the SIREN network with a softplus output is typically over-smoothed.
@@
Refer to captionRefer to caption
Figure 3: Visual results of the ablation studies: (a) Applying the absolute value to the output of the SIREN network. (b) Applying the softplus function to the output of the SIREN network. (c) Using uniform Eikonal constraints. (d) Removing normal alignment. (e) Replacing estimated normals with random vector. (f) Applying all components.
(a) DUDF
(b) CAP-UDF
@@
Figure 4: Visual comparisons with DUDF, CAP-UDF and LevelSetUDF on an indoor scene of the Stanford 3D Scene dataset featuring noise, an imperfect scan, and two car models of the ShapeNet-Cars dataset showcasing complex structures. Our method remains the open boundaries, such as, the bookshelves, inner structures of the vehicle, car window and car exhaust vent.
Normal alignment
To evaluate the contribution of normal alignment to reconstruct geometric details, we compare our method with and without the normal alignment loss. In addition, we test the robustness of our method when the estimated normals are unreliable. Figure 3 and Table 4 show the effects of the normal alignment loss. Interestingly, we observe that even fully random normals could also improve the results significantly. While the normals are unreliable, the $\mathcal{L}_{normal}$ also locates the local minimal distance surface between the two sample points $\mathbf{q}^{1}_{i}$ and $\mathbf{q}^{2}_{i}$. The is also beneficial to minimize the error.
CD-mean CD-median $F1^{0.005}$ $F1^{0.0025}$
SIREN+Abs. 5.27 3.07 87.65 77.87
SIREN+softplus 2.91 2.71 99.79 91.08
w/o weighted Eikonal 2.87 2.54 99.26 90.98
w/o normal alignment 3.20 2.97 99.36 88.68
random normal alignment 2.72 2.53 99.87 92.12
DEUDF 2.69 2.51 99.95 93.52
Table 4: Ablation studies on the model “Stonewall” (Figure 3) of the Stanford 3D Scene Dataset.
Weighted Eikonal
We explore the effects of different configurations of the Eikonal loss. Specifically, we compare our method using a standard Eikonal loss that is applied uniformly to all sample points with our adaptively weighted Eikonal loss. We observe that the standard Eikonal loss results in learned UDFs with lower accuracy near the zero level sets, leading to numerous small holes in the extracted meshes. In contrast, our adaptively weighted Eikonal loss more effectively addresses the vanishing gradient problem, and stabilizes the learning process, thereby yielding meshes with higher quality. In Figure 3 and Table 4, due to the uniform Eikonal constraint that impedes the UDF learning process at the zero level set, slightly higher distance values are observed around the target surface, which can result in holes in the extracted mesh, even though a larger DCUDF (Hou et al. 2023) threshold has be applied in this example.
Conclusions and Limitations
This paper presents an improved UDF learning method for high fidelity 3D surface reconstruction. The method integrates novel UDF representation, normal alignment, adaptively weighted Eikonal constraint and SIREN network to learn more accurate UDFs. Our DEUDF can not only lean geometry details but also keep boundaries, thereby maintaining better topology. Extensive experiments illustrate that our method produces low Chamfer distances and better topology outperforming state-of-the-art methods.
Our method is primarily designed for dense, evenly sampled point clouds, as sparse point clouds often fail to capture fine geometric details. If the input point cloud is highly uneven, in some sparse locations, the learned unsigned distance values could be high resulting in small holes in the reconstructed mesh. In the future, we aim to extend DEUDF to handle point clouds with non-uniform densities, where points are sparse in smooth areas and dense in regions with fine details.
Acknowledgments
This work was supported in part by the National Key R&D Program of China under Grant 2023YFB3002901, the Basic Research Project of ISCAS under Grant ISCAS-JCMS-202303, the Major Research Project of ISCAS under Grant ISCAS-ZD-202401, the Ministry of Education, Singapore, under its Academic Research Fund Grants (MOE-T2EP20220-0005 & RT19/22), and an OPPO gift fund.
Supplementary Material for
Details Enhancement in Unsigned Distance Field Learning for High-fidelity 3D Surface Reconstruction
Appendix
We present additional comparisons with CAP-UDF (Zhou et al. 2022) and LevelSetUDF (Zhou et al. 2023) by their original implementations as illustrated in Figure 6 and detailed in Table 5. The accuracy is lower than using DCUDF (Hou et al. 2023) for surface extraction. Our method produces higher quality results, by improved geometric detail and smoother shape boundaries. In Table 6, we show detailed results about the zero deviation from the zero level set. In Figure 5, we show visual comparisons for resilience to noisy input point clouds. In Figure 7 and 8, we show more results for models with details and models with complex topology and internal structures. The surfaces are all extracted by DCUDF except DUDF (Fainstein, Siless, and Iarussi 2024) by the original method. Our method outperforms DUDF, CAP-UDF and LevelSetUDF in terms of accuracy and topology.
@@ Acknowledgments
Table 5: Quantitative comparisons with the original results of CAP-UDF and LevelSetUDF. Chamfer distances are measured in the unit of ×$\times 10^{-3}$.
Stanford 3D Scene Stanford 3D Scan
Method Burghers Copyroom Lounge Stonewall Totempole Asian dragon Camera Dragon Dragon warrior Dragon wing Statue ramesses Thai statue Vase lion
DUDF 26.26 27.26 26.03 25.87 25.68 25.18 25.90 24.93 26.05 25.86 26.42 26.27 26.46
CAP-UDF 3.94 6.64 4.81 2.66 2.19 3.39 0.26 0.92 2.52 2.38 1.54 2.31 0.74
LeverSetUDF 1.98 3.63 3.06 0.83 1.12 1.24 0.31 0.49 0.80 0.72 1.90 1.50 0.45
Ours 1.43 1.61 1.48 0.92 0.99 1.02 0.94 1.48 2.22 1.92 0.53 1.69 1.26
Table 6: Quantitative results for deviations from zero. We input vertices from reconstructed mesh into learned UDF, then calculate the average output value, which means average deviation from zero. Our method performs better in average, and is more stable than other methods.
(a) CAP-UDF
(b) LevelSetUDF
(c) DUDF
(d) ours-30Freq
Figure 5: Visual results for reconstruction on noisy point clouds. We separately add $N(0,0.0025)$, $N(0,0.005)$ Gaussian noise to input point cloud. Specially, we changed our method to SIREN with frequency 30 for better performance.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(a) CAP-UDF
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(b) DUDF
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(c) LevelSetUDF
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(d) Ours
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
(e) GT
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 8: More visual comparisons with DUDF, CAPUDF and LevelSetUDF on car models from the ShapeNet-Cars dataset showcasing better reconstruction for complex internal structures.